OpenStudy (legends):

What is a perfect square? Is 0 a perfect square?

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (alexandervonhumboldt2):

0*0=0 sqrt0=0

Nnesha (nnesha):

perfect square when u take square root of number and get perfect number(not decimal) for example \[\sqrt{9} = 3\]

OpenStudy (alexandervonhumboldt2):

a perfect square is when the square of this number is an intenger

OpenStudy (legends):

So a perfect square is like square root??

OpenStudy (alexandervonhumboldt2):

yeah but the square root must be an intenger

OpenStudy (anonymous):

is \[\frac{9}{4}\] a prefect square ?

OpenStudy (anonymous):

how about \(1.44\) is that one?

OpenStudy (legends):

No

OpenStudy (anonymous):

ok if you say so

OpenStudy (legends):

?

OpenStudy (anonymous):

how about \(x^2+2x+1\) is that a perfect square?

OpenStudy (anonymous):

if you look up a definitinon you will see it is a square of a rational number, or a polynomial so since \(\frac{9}{4}\) is the square of \(\frac{3}{2}\) it is a prefect square

OpenStudy (anonymous):

likewise \[(1.2)^2=1.44\] making \[1.44\] a prefect square

OpenStudy (anonymous):

as well as \((x+1)^2=x^2+2x+1\)